OpenStudy (darkbluechocobo):

Solve the following quadratic equation: -2x^2 + x − 2 = 0

OpenStudy (darkbluechocobo):

Is the answer \[\left\{ 1/4 +/- \frac{ \sqrt{15} }{ 4i } \right\}\]